OpenStudy (anonymous):

Solve x2 = 14x – 49.

OpenStudy (anonymous):

x^2 - 14x + 49 = (x-7)^2

OpenStudy (anonymous):

thanks soo much but i need to know how to get it

OpenStudy (anonymous):

a^2 - 2ab + b^2 = (a-b)^2

OpenStudy (across):

moneybird factored that expression out by noticing that -7 + -7 = -14 and -7 * -7 = 49.

OpenStudy (anonymous):

...........i'm really confused can you clarify the work pleaseee